# JunoBench: A Benchmark Dataset of Crashes in Python Machine Learning Jupyter Notebooks

### v1

@@ Abstract.
Jupyter notebooks are widely used for machine learning (ML) prototyping. Yet few debugging tools are designed for ML code in notebooks, potentially due to the lack of benchmarks. We introduce JunoBench, the first benchmark dataset of real-world crashes in Python-based ML notebooks. JunoBench has 111 curated and reproducible crashes from public Kaggle notebooks, each paired with a verifiable fix, ranging over popular ML libraries, including TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, and NumPy, as well as notebook-specific out-of-order execution issue. To support reproducibility and ease of use, JunoBench offers a unified execution environment where crashes and fixes can be reliably reproduced. By providing realistic crashes and their resolutions, JunoBench facilitates bug detection, localization, and repair tailored to the interactive and iterative nature of notebook-based ML development.
@@ 1. Introduction
To fill this gap, we present JunoBench, the first benchmark of crashes in real-world Python ML notebooks. Based on our previous empirical study (Wang et al., 2025c), JunoBench (available at Hugging Face (Wang et al., 2025a)) supports the evaluation of debugging tools for notebook-based ML development by providing:
@@ 2.2. Benchmark inclusion
Manual reproducing and fixing procedure
@@
Benchmark balancing across ML libraries
We began with 174 labeled crash candidates and successfully reproduced and fixed 74 of them. To balance ML library coverage, we continued with the additionally sampled 500 unlabeled ones. We first applied our previously introduced heuristic tool to automatically infer the library cause of each crash, based on which we prioritized our reproduction and fix efforts.To avoid overrepresentation of any single library, we monitored the distribution as we processed. Once a library became sufficiently represented, we halted further reproduction efforts for crashes associated with that library and focused on others. From this unlabeled set, we evaluated 93 additional crashes and successfully reproduced and fixed 37. Unreproducible crashes were primarily due to missing or inaccessible datasets.
@@
Most datasets permit research use. In four cases where the original datasets (all numeric) are under a private or restrictive license, we replaced them with synthetic datasets by replicating the original dataset’s schema and generating random values. Each input dataset in JunoBench (Wang et al., 2025a) is accompanied by a metadata file that documents its title, source URL, license, and modifications applied (e.g., downsampling, synthetic substitution).
@@ 2.3. Benchmark usability
To improve usability, we provide a unified execution environment for all notebooks in JunoBench. We built a Docker image based on the official Kaggle notebook platform (Kaggle, 2025), extended it with additional dependencies required by the notebooks. This environment ensures all cases can be reproduced consistently with a one-time setup. The full dependency specification is provided in our benchmark repository (Wang et al., 2025a).

### v2

@@ Abstract.
Jupyter notebooks are widely used for machine learning (ML) prototyping. Yet few debugging tools are designed for ML code in notebooks, potentially due to the lack of benchmarks. We introduce JunoBench, the first benchmark dataset of real-world crashes in Python-based ML notebooks. JunoBench has 111 curated and reproducible crashes from public Kaggle notebooks, each paired with a verifiable fix, ranging over popular ML libraries, including TensorFlow/Keras, PyTorch, Scikit-learn, Pandas, and NumPy, as well as notebook-specific out-of-order execution issue. To support reproducibility and ease of use, JunoBench offers a unified execution environment where all crashes can be reliably reproduced. By providing realistic crashes and their resolutions, JunoBench facilitates bug detection, localization, diagnosis, and repair tailored to the interactive and iterative nature of notebook-based ML development.
@@ 1. Introduction
To fill this gap, we present JunoBench, the first benchmark of crashes in real-world Python ML notebooks. Based on our previous empirical study (Wang et al., 2025c), JunoBench supports the evaluation of debugging tools for notebook-based ML development by providing:
@@
  • Ground truth crash diagnosis labels for developing and evaluating crash identification tools.
@@
JunoBench is available on Hugging Face (Wang et al., 2025a), and all artifacts used to construct the benchmark are provided on GitHub (Wang et al., 2025b).
@@ 2.2. Benchmark inclusion
Manual reproducing and fixing
@@
To construct a balanced benchmark across ML libraries, we began with 174 labeled crash candidates and successfully reproduced and fixed 74 of them. Then we continued with the additionally sampled 500 unlabeled ones. We first applied our previously introduced heuristic tool to automatically infer the library cause of each crash, based on which we prioritized our reproduction and fix efforts.To avoid overrepresentation of any single library, we monitored the distribution as we processed. Once a library became sufficiently represented, we halted further reproduction efforts for crashes associated with that library and focused on others. From this unlabeled set, we evaluated 93 additional crashes and successfully reproduced and fixed 37. Unreproducible crashes were primarily due to missing or inaccessible datasets.
@@
Crash detection and diagnosis annotation
To support the development and evaluation of crash identification tools, we provide ground truth labels to standardize validation. Specifically, we define: (1) the ground truth crash detection label, assigned at the notebook level, where true denotes a reproduced version and false denotes a fixed version; and (2) the ground truth crash diagnosis label describing the cause (i.e., localization with a natural language explanation) of each crash in the reproduced versions. For all 111 crashes, the diagnosis labels were first created by a primary annotator based on three sources of evidence: (i) the crashing code cells, (ii) the corresponding error outputs (including error messages and tracebacks), and (iii) the corresponding fixed versions revealing how crashes were resolved. A second annotator then independently validated the diagnosis labels using the same information. In cases of disagreement, the two annotators conducted discussions until full consensus was achieved.
@@
Most datasets permit research use. In four cases where the original datasets (all numeric) are under a private or restrictive license, we replaced them with synthetic datasets by replicating the original dataset’s schema and generating random values. Each input dataset in our benchmark is accompanied by a metadata file that documents its title, source URL, license, and modifications applied (e.g., downsampling, synthetic substitution).
@@ 2.3. Benchmark usability
To improve usability, we provide a unified execution environment for all notebooks in JunoBench. We built a Docker image based on the official Kaggle notebook platform (Kaggle, 2025), and then extended it with additional dependencies required by the notebooks. This environment ensures all cases can be reproduced consistently with a one-time setup.